OpenStudy (anonymous):

Felix purchased a boat for $9,755. He made a down payment of $875. He applied for a six-year installment loan with an interest rate of 10.4% in the amount of $8,880. What is the total cost of the boat after six years? $9,803.52 $11,974.32 $12,849.32 $10,769.52

OpenStudy (phi):

I can answer this question using two formulas \[ FV= PV(1+i)^n \] Future Value equals Present Value times 1+interest raised to the nth power and \[ FV= pymt \cdot \frac{(1+i)^n-1}{i} \] This says Future Value is the payment times the complicated mess... I would figure out the Future Value, and then use that to find the payment The payment times the number of months paid is how much you pay for the boat (and add in the down payment)

OpenStudy (phi):

I assume a six-year installment plan has monthly payments for 72 months (6 years * 12 months/year) interest (as used in the formula) is 0.104/12 (interest per month) n is 72 (months) PV present value is 8880 (amount of the loan) FV= 8880.0 * (1+ 0.104/12)^72 type 8880.0 * (1+ 0.104/12)^72 = into the google search window you should get 16528.9449

OpenStudy (phi):

Can you find the monthly payment? If you do, round it to the nearest penny. Then multiply it by 72 to get the amount paid. The final answer is that amount plus the down payment of 875.00

OpenStudy (anonymous):

your wrong. where do you see one of those answers?